A uniform ladder rests against a smooth wall so that it makes an angle of 600 with the ground. The ladder is 10.0 m long and weighs 150 N. How far can a 250 N child go up the ladder before it gets slips?

The coefficient of friction between the ladder and the ground is 0.4.


1
Expert's answer
2021-02-15T00:54:24-0500

πœ‡1 = 0.400 βˆ’ coefficient of friction between the ladder and the ground;

πœ‡2 = 0.450 βˆ’ coefficient of friction between the ladder and the wall;

𝛼 = 60Β° βˆ’ angle which ladder makes with the ground;

𝑁1 βˆ’ reaction force from the ground;

𝑁2 βˆ’ reaction force from the wall;

𝑃1 = 150𝑁 βˆ’ weight of the ladder;

𝑃2 = 250𝑁 βˆ’ weight of the child;

𝐿 = 10.0m βˆ’ length of the ladder.


We will consider the extreme case when the person is standing at a maximum distance d from the beginning of the ladder.

Newton's second law for the ladder (the first law of equilibrium):

"\vec{F_{fr1}}+\vec{F_{fr2}}+\vec{P_1}+\vec{P_2}+\vec{N_1}+\vec{N_2}=\vec{0},"

Projection of the law on the X-axis:

"N_2-F_{fr1}=0," (1)

projection of the law on the X-axis:

"N_1+F_{fr2}-P_1-P_2=0." (2)

Law of dry friction:

"F_{fr1}=\mu_1N_1" (3)

"F_{fr2}=\mu_2N_2" (4)

(3) and (4) in (1) and (2):

(3) β†’ (1): Β "N_2-\mu_1N1=0"

"N_1=\frac{N_2}{\mu_1}" (5)

(4) β†’ (2): "N_1+\mu_2N_2-P_1-P_2=0"

(5) β†’ (2):

"\frac{N_2}{\mu_1}+\mu_2N_2-P_1-P_2=0" ,

"N_2=\frac{\mu_1(P_1+P_2)}{1+\mu_1\mu_2}=135.6~\text{N},"

"F_{fr2}=\mu_2N_2=61.02~\text{N}."

The second law of equilibrium:

"M_{P_1}+M_{P_2}+M_{F_{fr2}}+M_{N_2}=0," (6)

("M_{N_1}=M_{F_{fr1}}=0" π‘π‘’π‘π‘Žπ‘’π‘ π‘’ π‘šπ‘œπ‘šπ‘’π‘›π‘‘ π‘Žπ‘Ÿπ‘š π‘œπ‘“ π‘‘β„Žπ‘–π‘  π‘“π‘œπ‘Ÿπ‘π‘’π‘  𝑖𝑠 π‘§π‘’π‘Ÿπ‘œ),

"M_{P_1}=-P_1\cdot \frac L2 cos\alpha" (π‘šπ‘–π‘›π‘’π‘  𝑠𝑖𝑔𝑛 π‘π‘’π‘π‘Žπ‘’π‘ π‘’ π‘œπ‘“ π‘‘β„Žπ‘’ π‘‘π‘–π‘Ÿπ‘’π‘π‘‘π‘–π‘œπ‘› π‘œπ‘“ π‘“π‘œπ‘Ÿπ‘π‘’)

"M_{P_2}=-P_2\cdot dcos\alpha,"

"M_{F_{fr2}}=F_{fr2}\cdot Lcos\alpha,"

"M_{N_2}=N_2\cdot Lsin\alpha,"

β†’ (6):

"F_{fr2}\cdot Lcos\alpha+N_2\cdot Lsin\alpha-P_2\cdot dcos\alpha-P_1\cdot \frac L2 cos \alpha=0,"

"d=\frac{1}{P_2 cos\alpha}\cdot(F_{fr2}\cdot Lcos\alpha+N_2\cdot Lsin\alpha-P_1\cdot \frac L2 cos \alpha)=8.84~\text{m}."